# GridPE: Unifying Positional Encoding in Transformers with a Grid Cell-Inspired Framework

### v1

@@ 4.1 Property of Distance Decay
4.2 Performance of GridPE on Visual Task
To empirically validate the capability of GridPE strategies to encode positional information, we draw on previous research [30, 31] and design a image classification task to evaluate neural network architectures with different positional encodings. We select the Pyramid Vision Transformer (PVT) as the base architecture [20]. PVT, a classic variant of the Vision Transformer, introduces a pyramid structure that reduces input resolution and computation cost, thereby enhancing task performance.
We utilize the ImageNet-100 dataset to design the classification task. We evenly select ten classes [^1] These classes are: American robin, American lobster, Saluki, poodle, hare, car wheel, honeycomb, mousetrap, safety pin, vacuum cleaner, and for each class, we pick 250 images. Among them, 50 evenly chosen images are partitioned as the validation set, while the other 200 are used for training. We compare seven models with different positional encoding choices. All positional encoding approaches are applied to the base attention-based structure. PVT-o refers to the original PVT architecture with no explicit two-dimensional positional encoding. PVT-abs refers to applying classical and sinusoidal two-dimensional positional encoding to image pixels [28]. CPVT is a variant of PVT with conditional positional encoding [31]. GridPVT-rotate, GridPVT-merge, GridPVT-complex, and GridPVT-deep correspond to the models enhanced with the four aforementioned GridPE encoding methods, respectively. We train the neural networks on a device with a Ubuntu 22.04 operation system, an Intel Xeon Platinum 8352V CPU, and an NVIDIA RTX 4090 GPU. The model are trained on the training set for 200 epochs, with AdamW optimizer.
The performance of all trained models on the validation set is shown in Table 1. We can see that our proposed GridPVT model supported by GridPE enhances classification performance to some extent. Although GridPVTs display similar performance to CPVT in the top-1 prediction task, for top-5 classification, the accuracy of GridPVT models generally surpasses that of the other three models. Among the four GridPVTs, the model with positional code merging and neural network projection achieves better accuracy scores than the others.
@@ 4.2 Performance of GridPE on Visual Task
Model Name Acc@1 (%) Acc@5 (%) Loss
PVT-o 58.6 92.8 1.467
CPVT 60.6 91.8 1.498
PVT-abs 58.6 92.8 1.483
GridPVT-rotate 54.2 93.4 1.511
GridPVT-merge 56.0 94.8 1.507
GridPVT-complex 56.4 93.6 1.514
GridPVT-deep 59.2 94.2 1.445
5 Conclusion
Positional encoding is crucial for artificial intelligence as it elucidates how information is effectively organized within neural networks. This study introduces GridPE, a novel positional encoding framework inspired by the neural strategies of grid cells in the mammalian brain, to enhance transformer architectures for high-dimensional data process. GridPE models the position encoding inspired by VCO theory of grid cells, which not only captures the absolute positioning within high-dimensional spaces but also preserves the relative positioning essential for tasks that require an understanding of spatial relationships. The image classification experiment showcase the performance of the model enhanced by GridPE, with classification accuracy not worse than state-of-art positional encoding apporaches. The versatility and scalability of GridPE, evident from its effective application across high dimensions of input data , pave the way for future exploration in natural language processing, robotics, and spatial cognition. By framing spatial understanding as an inverse Fourier transform, GridPE not only enhances artificial intelligence with neuroscience principles but also offers a sophisticated, biologically inspired framework for positional encoding.

### v2

@@ Introduction
Performance of GridPE on Visual Task
To empirically validate the capability of GridPE strategies in encoding positional information, we draw on previous research (Chu et al. 2021, 2023) and design an image classification task to evaluate neural network architectures with different positional encodings. We select the Pyramid Vision Transformer (PVT) (Wang et al. 2021) and Vision Transformer (ViT) (Dosovitskiy et al. 2020) as the base architectures.
Baseline model descriptions.
PVT introduces a pyramid structure into the transformer architecture, which reduces input resolution and computational cost while enhancing performance in dense prediction tasks by effectively capturing multi-scale features. ViT, on the other hand, revolutionizes image classification by applying the transformer model to fixed-size image patches, leveraging global self-attention to model relationships across the entire image.
Dataset.
We utilize the ImageNet-100 dataset, a subset of the widely recognized ImageNet dataset, to design our classification task. With its diverse and complex imagery, ImageNet-100 serves as an ideal benchmark for evaluating the effectiveness of GridPE in encoding positional information, rigorously testing the model’s ability to handle high-dimensional and varied visual inputs. For our task, we partition 20% of the images into a validation set, while the remaining 80% are used for training, ensuring robust model evaluation and effective training during the experiment.
Implementation details.
We train the neural networks on a system with an NVIDIA RTX 4090 GPU using the PyTorch framework. Models are trained for 200 epochs with the AdamW optimizer and a batch size of 32. The ‘ReduceLROnPlateau‘ scheduler is employed to halve the learning rate if validation accuracy does not improve for 10 consecutive epochs, helping to fine-tune the model and prevent overfitting. PVT models use an image size of 384, while ViT models use 224. These settings ensure robust training and evaluation on the ImageNet-100 dataset.
Results and Analysis
We assess the effectiveness of GridPE by comparing models with different positional encoding strategies applied to two core attention-based architectures: PVT and ViT. For PVT, we evaluate the original version without explicit two-dimensional positional encoding (PVT-o), a variant using classical sinusoidal encoding (PVT-abs) (Vaswani et al. 2017), and a model incorporating conditional positional encoding (CPVT) (Chu et al. 2023). Additionally, we introduce four GridPE-enhanced versions—GridPVT-rotate, GridPVT-complex, GridPVT-deep, and GridPVT-merge—each reflecting different grid cell-inspired positional encoding strategies. We also apply these GridPE variants to the ViT architecture to examine their impact on a simpler and less parameter-intensive model. By comparing these GridPE-enhanced versions with the baseline ViT model, we aim to elucidate the influence of GridPE across models with varying architectural complexities.
Model Name Acc@1 (%) Acc@5 (%) Loss
PVT-o 84.1 96.2 1.467
CPVT 87.1 95.8 1.498
PVT-abs 85.3 96.6 1.483
GridPVT-rotate 82.3 95.7 1.511
GridPVT-complex 83.0 95.9 1.507
GridPVT-deep 84.6 97.4 1.445
GridPVT-merge 82.9 95.8 1.514
ViT 76.1 93.0 0.8234
GridViT-rotate 77.6 91.5 0.9832
GridViT-complex 78.8 93.7 0.8691
GridViT-deep 81.5 94.2 0.4676
GridViT-merge 79.0 94.0 1.083
@@
Table 1 presents the comparative results, showing how different positional encoding strategies impact the performance of PVT and ViT models. Among these PVT models, CPVT achieves the highest top-1 accuracy (87.1%), while GridPVT-deep excels with the best top-5 accuracy (97.4%) and the lowest loss (1.445). The convergence curves in Figure 4(a) reveal that PVT models generally maintain stable and efficient convergence regardless of the positional encoding used, likely due to their complex architecture and larger parameter count. This complexity allows the PVT models to capture intricate spatial relationships, minimizing the impact of different positional encodings on their convergence speed and overall performance.
For the ViT architecture, the application of GridPE variants results in noticeable improvements in performance, as seen in Table 1. GridViT-deep emerges as the most successful variant, achieving the highest top-1 accuracy (81.5%) and the lowest loss (0.4676), representing significant gains over the baseline ViT model. The validation accuracy curves in Figure 4(b) highlight that ViT models, which are simpler and have fewer parameters compared to PVT models, are more sensitive to changes in positional encoding. While other variants such as GridViT-complex and GridViT-merge also outperform the baseline ViT model, they exhibit higher losses, suggesting that while GridPE can enhance accuracy, it may require further tuning to optimize efficiency and convergence. This underscores the importance of carefully matching positional encoding strategies with the model architecture to achieve the best performance outcomes.
(a) Validation Accuracy over Epochs for Multiple Variants of PVT Models.
(b) Validation Accuracy over Epochs for Multiple ViT Models.
Figure 4: Comparison of Validation Accuracy over Epochs for ViT and PVT Models.
Conclusion
Positional encoding is essential for structuring information in neural networks, with its generality suggesting a unified algorithmic strategy that applies across visual, auditory, and linguistic domains. This study introduces GridPE, a unified framework inspired by grid cells in the mammalian brain, to enhance transformer architectures for high-dimensional data processing. GridPE captures both absolute and relative positioning, crucial for tasks involving spatial relationships. Experiments in image classification demonstrate that GridPE achieves competitive accuracy compared to state-of-the-art approaches. Its versatility and scalability across different data domains suggest promising applications in natural language processing, robotics, and spatial cognition, blending neuroscience principles with artificial intelligence.